OpenStudy (anonymous):

Suppose that the graph of a function f is known. Explain how the graph of y = 4f(x) differs from the graph of y = f(4x).

pooja195 (pooja195):

y = 4f(x) stretches vertically the graph of y = f(x) by a factor of 4. y = f(4x) compresses horizontally the graph of y = f(x) by a factor of 4.